# xxxxx, xxxxx, xxxx

### v3

@@ 9 Conclusions: Why Nature Does Not Like Zombies
We have described a multilayered formalism illustrating how biological self-organising systems become phenomenally conscious when they construct a 1ST order self. A human lacking a 1ST order self could not perform causal reasoning needed to adapt as humans evidently can [pearl2018].
As previous research pointed out [bennett2024a], the computer metaphor with hardware and software is a simplification. Rather, software is a state of hardware. Nature privileges efficiency over abstract simplification. Compare the vast quantities of both training data and energy required by a large language model, to the small quantities humans need to solve a problem. Biological systems are more efficient, because they are adaptive at a every level [levin2024, bennett2024c]. Hence, rather than trying to explain the mind in the abstract like software, we started at the level of the embodied organism.
@@
Furthermore, the conceivability of zombies depends on the ability to duplicate disposition but not phenomenal experience. Carruth argued that if qualia are dispositional, then that is impossible [carruth2016]. Our arguments align with Carruth’s. A causal identity is inherently dispositional, as is embodiment.
@@
Appendix A Appendix
For convenience of reference, we have placed all definitions here. We have aimed to make the gist of the paper understandable without the math, hence the reader can either skip this section or refer back to it if needed. Many of the definitions have been adapted from a variety of preceding work [bennett2024a, bennett2023b, bennett2023c, bennett2023d, bennett2024b, bennett2024c]. They are referred to in the body of the paper when they become relevant, in the order in which they appear here.
@@ Appendix A Appendix
(optimal proxy) There is no proxy more sample efficient than weakness. The weakness proxy formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [bennett2023b]).
(intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy. Humans typically learn from “examples”. An example of a task is a correct output and input. A collection of examples is a child task, so “learning” is an attempt to generalise from a child to one of its parents. The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy. The most sample efficient proxy is weakness [bennett2023b, prop. 1, 2], which is why we’re using it here.
@@
(intuitive summary) An intervention is action undertake an organism or other agency, in the sense described by Pearl [pearl2018]. Intuitively, if “$int$” and “$obs$” are events which have happened, then we say that $int$ has caused $obs$ if $obs$ would not have happened in the absence of $int$ (counterfactual).
@@
A.1 Proofs
Versions of or alternatives to proofs 1-3 were published in [bennett2023b]. Further details on these updated versions are available in on GitHub, along with the supporting experiments [bennett2024appendices].
Theorem 1 (sufficiency).
@@ A.1 Proofs
Theorem 3 (simplicity sub-optimality).
@@
    From [bennett2023b, prop. 3] we have proof that weakness is the optimal choice of proxy to maximise the probability of generalisation from child to parent is the weakest policy. It follows that $\mathfrak{a}$ will generalise from $\mathfrak{h}_{<t_{\mathfrak{a}}}$ to $\mu_{\mathfrak{a}}$ given the smallest history of interventions with which it is possible to do so (meaning the smallest possible ostensive definition, or cardinality $\lvert D_{\alpha}\rvert$).

### v4

@@ 9 Conclusions: Why Nature Does Not Like Zombies
We have described a multilayered formalism illustrating how biological self-organising systems become phenomenally conscious when they construct a 1ST order self. A human lacking a 1ST order self could not perform causal reasoning needed to adapt as humans evidently can [56].
As previous research pointed out [31], the computer metaphor with hardware and software is a simplification. Rather, software is a state of hardware. Nature privileges efficiency over abstract simplification. Compare the vast quantities of both training data and energy required by a large language model, to the small quantities humans need to solve a problem. Biological systems are more efficient, because they are adaptive at a every level [23, 38]. Hence, rather than trying to explain the mind in the abstract like software, we started at the level of the embodied organism.
@@
Furthermore, the conceivability of zombies depends on the ability to duplicate disposition but not phenomenal experience. Carruth argued that if qualia are dispositional, then that is impossible [89]. Our arguments align with Carruth’s. A causal identity is inherently dispositional, as is embodiment.
@@
10 Appendix
For convenience of reference, we have placed all definitions here. We have aimed to make the gist of the paper understandable without the math, hence the reader can either skip this section or refer back to it if needed. Many of the definitions have been adapted from a variety of preceding work [31, 33, 34, 50, 36, 38]. They are referred to in the body of the paper when they become relevant, in the order in which they appear here.
@@ 10 Appendix
(optimal proxy) There is no proxy more sample efficient than weakness. The weakness proxy formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [33]).
(intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy. Humans typically learn from “examples”. An example of a task is a correct output and input. A collection of examples is a child task, so “learning” is an attempt to generalise from a child to one of its parents. The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy. The most sample efficient proxy is weakness [33, prop. 1, 2], which is why we’re using it here.
@@
(intuitive summary) An intervention is action undertake an organism or other agency, in the sense described by Pearl [56]. Intuitively, if “$int$” and “$obs$” are events which have happened, then we say that $int$ has caused $obs$ if $obs$ would not have happened in the absence of $int$ (counterfactual).
@@
10.1 Proofs
Versions of proofs 1-3 were published in proceedings [33]. Further details are available in on GitHub, along with the supporting experiments [35].
Proposition 1 (sufficiency).
@@ 10.1 Proofs
Proposition 3 (simplicity sub-optimality).
@@
    From [33, prop. 3] we have proof that weakness is the optimal choice of proxy to maximise the probability of generalisation from child to parent is the weakest policy. It follows that $\mathfrak{a}$ will generalise from $\mathfrak{h}_{<t_{\mathfrak{a}}}$ to $\mu_{\mathfrak{a}}$ given the smallest history of interventions with which it is possible to do so (meaning the smallest possible ostensive definition, or cardinality $\lvert D_{\alpha}\rvert$).